Point B is the midpoint of AC .
A (2, -1)
B (k, t)
C (2, 5)
What is the value of t?

Answer by jim_thompson5910(35256)

t is equal to the average of the y coordinates of points A and B


So t = (-1+5)/2 = 4/2 = 2


which means that t = 2